OpenStudy (anonymous):

What is the vertex of the graph of y = 2(x – 3)2 + 4? (4, 3) (4, –3) (3, 4) (–3, 4)

OpenStudy (anonymous):

Is there a picture of the graph?

OpenStudy (zzr0ck3r):

\(ax^2+bx+c\) has vertex at \(\frac{-b}{2a}\)

OpenStudy (zzr0ck3r):

and the point is given by \((\frac{-b}{2a},f(\frac{-b}{2a}))\)

OpenStudy (zzr0ck3r):

\(2(x-3)^2+4=2(x^2-6x+9)+4=2x^2-12x+18+4 = 2x^2-12x+22\) so the vertex is at \(x=\frac{12}{4}=3\) plug that into the polynomial and get f(3) = 4 so the vertex is at (3,4)

OpenStudy (zzr0ck3r):

*so the vertex is (3,4)

OpenStudy (anonymous):

thanks